On July 1, 2019, Carla Vista Co. purchased new equipment for $85,000. Its estimated useful life was 5 years with a $10,000 salvage value. On December 31, 2022, the company estimated that the equipment’s remaining useful life was 10 years, with a revised salvage value of $5,000.

Correct answer iconYour answer is correct.

Prepare the journal entry to record depreciation on December 31, 2019. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Account Titles and Explanation

Debit

Credit

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

eTextbook and Media

List of Accounts

Correct answer iconYour answer is correct.

Prepare the journal entry to record depreciation on December 31, 2020. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Account Titles and Explanation

Debit

Credit

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

eTextbook and Media

List of Accounts

Incorrect answer iconYour answer is incorrect.

Compute the revised annual depreciation on December 31, 2022.

Revised annual depreciation $enter the Revised annual depreciation in dollars

eTextbook and Media

List of Accounts

Partially correct answer iconYour answer is partially correct.

Prepare the journal entry to record depreciation on December 31, 2022. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Account Titles and Explanation

Debit

Credit

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

eTextbook and Media

List of Accounts

Incorrect answer iconYour answer is incorrect.

Compute the balance in Accumulated Depreciation—Equipment for this equipment after depreciation expense has been recorded on December 31, 2022.

Accumulated Depreciation—Equipment
Answer #1

a) Journal entry

Account title and explanation Debit Credit
Depreciation expense (85000-10000/5)*6/12 7500
Accumulated Depreciation—Equipment 7500

b) Journal entry

Account title and explanation Debit Credit
Depreciation expense (85000-10000/5)*6/12 15000
Accumulated Depreciation—Equipment 15000

c) Revised annual depreciation = (85000-37500-5000)/10 = 4250

d) Journal entry

Account title and explanation Debit Credit
Depreciation expense 4250
Accumulated Depreciation—Equipment 4250

e) Accumulated depreciation-equipment = 37500+4250 = 41750